oose the correct interpretation of the standard deviation below. A. The standard deviation represents the sum of the deviations from the mean. B. The standard deviation represents finding the deviation for each observation, squaring each deviation, and then adding them up. C. Since the standard deviation uses the square of the units of measurement for the original data, it is not easy to interpret. D. The standard deviation represents a typical distance of an observation from the mean. Suppose the 7 was incorrectly recorded and is supposed to be 70. Redo (a)-(b) with the correct data and describe the effect of this outle he new range is 70 days (Round to two decimal places as needed.)
A company decides to investigate the amount of sick leave taken by its employees. A sample of seven employees yields the following numbers of days of sick leave taken in the past year 0 140 0 72 D Use this information to answer parts a through c a. Find and interpret the range. The range is 7 days. Choose the correct interpretation of the range below. A. The number of days separating the fewest and most sick days taken is equal to the range. B. The range represents the average distance of an observation from the mean. C. The largest difference between the mean and any other value is equal to the range. D. The range gives the most useful value for measuring the spread of the data. b. Find and interpret the standard deviation s. s# 2.65 (Round to two decimal places as needed.)
Answer #1

Solution :- New Data: o 1400 70 2 {m=7} -o V s.d = {(xi -a let us calculate N-1 ř = 0+1+ 4 + otot 70+ 2 0 + 0+ 70 + 21 = 10-